(Intermediate Macroeconomics)

2.The aggregate supply function be ys=2000+P, and the aggregate demand function be yD=2400-P。Find the equilibrium point between supply and demand.
Answer #1

At equilibrium point, aggregate demand = aggregate supply.

yD = yS

2400-P = 2000+P

2400 - 2000 = P = P

400 = 2P

P = 400 / 2

P = 200 (Equilibrium Price)

Note: yD = yS = y (At equilibrium)

y = 2400 - P

y = 2400 - 200

y = 2200 (Equilibrium output)